OpenStudy (anonymous):

r=sin(theta) and r=1-sin(theta) give an algebraic solution to all points of intersection

OpenStudy (cwrw238):

sin(theta) = 1 - sin(theta) sin(theta) = 1/2 theta = 30 degrees for theta 0 - 360 solutions are 30, 150

OpenStudy (anonymous):

algebraic solution?